Compute the torque on a shaft transmitting 750 W of power while rotating at 183 rad/s. This is equivalent to a 1.0 hp, 4-pole electric motor operating at its rated speed of 1750rpm.
Answer #1

Given data:

Power, P= 750 W

Rotation, \omega= 183 rad/s.

Now torque, T (N-m) is calculated as per following procedure:

P=T\times \omega

T=

750 183 T=

T= 4.09836 N-m. .....................Answer

Calculate the torque of given electric motor

We have the following relation for power, P

2π NT P= 60

Where N= RPM

T is Torque in N-m

P =power in Watt.

Now 1 HP is equal to near about 750 W.

Putting all given values.

P×60 T=

750 × 60 2π × 1750 T=

T= 4.09836 N-m.....Answer.

So given answer is correct.